OpenStudy (anonymous):

Evaluate. w = 3, x = 6, y = 5, and z = 9 xy+wz

OpenStudy (anonymous):

w = 3, x = 6, y = 5, and z = 9 xy+wz (6x5)+(3x9)= 30+27 = 57